Solve the following equations:
2q = 6
Solve the following equations:
2q = 6

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Now,
We have to divide both sides of the equation by 2,
Then, we get,
= 2q/2 = 6/2
= q= 3

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows :
2q = 6
q= 6/2
q = 3
 
Hope it helps!
Thank you ✌️